# Markovian Flow Matching: Accelerating MCMC with Continuous Normalizing Flows

### v1

@@ 3.2 Adaptive MCMC with Flow Matching
The output of Algorithm 1 is a vector of parameters $\theta_{K}$ which defines a CNF $\smash{(\phi_{t}^{\theta_{K}})_{t\in[0,1]}}$. Under the assumption that the parameter estimate converges, that is, $\smash{\theta_{K}\rightarrow\theta^{*}_{\mathrm{global}}}$ as $K\rightarrow\infty$, where $\smash{\theta^{*}_{\mathrm{global}}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathcal{J}(\theta;\pi)}$ denotes the global minimizer of the CFM objective $\mathcal{J}(\theta;\pi)$ in (5), this CNF is guaranteed to generate a probability path $(p_{t}^{\theta})_{t\in[0,1]}$ which transports samples from the reference $p_{0}$ to the true target $\pi$ [e.g., 42].
In practice, the objective $\smash{\mathcal{J}(\theta;\pi)}$ is highly non-convex, and thus it is not possible to establish a convergence result of this type without imposing unreasonably strong assumptions on the vector field $\smash{(v_{t}^{\theta})_{t\in[0,1]}}$. This being said, it is reasonable to ask whether $\theta_{K}$ converges to a local optimum of the CFM objective. We now answer this question in the affirmative. In particular, under mild regularity conditions, Proposition 3.1 guarantees that $\theta_{K}\rightarrow\theta^{*}$ almost surely as $K\rightarrow\infty$, where $\theta^{*}$ denotes a local minimum of the CFM objective. The proof of this result, which relies on a classical result in [5, Theorem 3.17], is provided in Appendix B.
@@ 4 Related work
In recent years, a number of works have proposed algorithms which combine MCMC techniques with NFs; see, e.g., [25, 1] for recent surveys. Broadly speaking, these algorithms fall into two categories. NeutraMCMC methods leverage NFs as reparameterization maps which simplify the geometry of the target distribution, before running (local) MCMC samplers in the latent space. This technique was pioneered in the seminal paper [56], and since been investigated in a number of different works [e.g., 30, 51, 55, 41, 79, 76, 65, 11]. Flow MCMC methods, meanwhile, utilise the pushforward of the base distribution through the NF as an (independent) proposal within an MCMC scheme. This approach was first studied by [2], and further extended in [54, 3, 28]. More recently, [21, 64] have introduced adaptive MCMC schemes which combine local MCMC samplers (e.g., MALA or HMC), with a non-local, flow-informed proposal (IMH or i-SIR); see also [32]. Our algorithm combines aspects of both neutraMCMC and flow MCMC methods and, unlike any existing approach, make use of a CNF (as opposed to a discrete NF), by leveraging the conditional flow matching objective.
The use of NFs within other Monte Carlo algorithms has also been the subject of recent interest. For example, [4, 47] consider augmenting SMC with NFs, while [16, 49] use NFs (or diffusion models) within AIS. Finally, although not directly comparable to our own approach, several recent works have proposed to use (controlled) diffusion processes to sample from unnormalized probability distributions. Such methods include [80], who introduce the path integral sampler, [71], who propose the denoising diffusion sampler, and [77] who introduce generative flow samplers. Other relevant contributions in this direction include [69, 6, 60, 72, 71, 73, 70, 74, 57].
@@ 5 Experiments
In this section, we evaluate the performance of MFM (Algorithm 1) on two synthetic and two real data examples. Our method is benchmarked against three relevant methods. The Denoising Diffusion Sampler [DDS; 71] is a VI method which approximates the reversed diffusion process from a reference distribution to an extended target distribution by minimizing the KL divergence. Adaptive Monte Carlo with Normalizing Flows [NF-MCMC; 21] is an augmented MCMC scheme which uses a mixture of MALA and adaptive transition kernels learned using discrete NFs. Finally, Flow Annealed Importance Sampling Bootstrap [FAB; 49] is an augmented AIS scheme minimizing the mass-covering $\alpha$-divergence with $\alpha=2$.
@@
$v_{t}^{\theta}(x)=\text{NN}(x,t;\theta_{1})+\text{NN}(t;\theta_{2})\times\nabla\log\pi(x),$ (17)
where the neural networks are standard MLPs with 2 hidden layers, using a Fourier feature augmentation for $t$ [67]. This architecture is also used by DDS [71, Section 4]. Meanwhile, FAB and NF-MCMC use rational quadratic splines [18]. Flows are trained using Adam [37] with a linear decay schedule terminating at $\varepsilon_{K}=0$. We report results for all methods averaged over 10 independent runs with varying random seeds. The code to reproduce the experiments is provided at albcab/mfm.
@@
Our first example is a mixture of four Gaussians, evenly spaced and equally weighted, in two-dimensional space. The four mixture components have means $(8,8)$, $(-8,8)$, $(8,-8)$, $(-8,-8)$, and all have identity covariance. This ensures that the modes are sufficiently separated to mean that jumping between modes requires trajectories over sets with close to null probability. Given the synthetic nature of the problem, we can measure approximation quality using the Maximum Mean Discrepancy (MMD) [e.g., 26]; see Appendix C.1 for details. We can also include, as a benchmark, the results for an approximation learned using FM with true target samples. Diagnostics for all models are presented in Table 1, and learned flow samples in Fig. 1. Further algorithmic details and results are provided in Appendix C.2.
@@
In this experiment, only our method (Fig. 1(a)) and DDS (Fig. 1(c)) learn the fully separated modes, reflecting the greater expressivity of CNFs in comparison to the discrete NFs used in, e.g., NF-MCMC (Fig. 1(d)). It is worth noting that DDS provides a closer approximation to the real target than MFM and, notably, even FM trained using true target samples (top row). Given that both methods use the same network architecture but a different learning objective, this suggests a potential limitation with the FM objective, at least when using this network architecture. This being said, MFM is notably more efficient than DDS (as well as the other methods) in terms of total computation time. While this is not a critical consideration in this synthetic, low-dimensional setting, it is a significant advantage of MFM in higher-dimensional settings involving real data (e.g., Section 5.3 and Section 5.4).
@@
The second experiment is a mixture of bivariate Gaussians with 16 mixture components. This is a modification of the 4-mode example, with contrasting qualities that illustrate other characteristics of each of the presented methods. In this case, the modes are evenly distributed on $[-16,16]^{2}$, with random log-normal variances. The number of modes reduces the size of sets of (near) null probability between the modes, making jumping between them easier. To increase the difficulty of this model, all methods are initialized on a concentrated region of the sampling space. Diagnostics are presented in Table 1 and learned flow samples in Fig. 2. Further details are provided in Appendix C.3.
In this example, DDS collapses to the modes closest to the initial positions while our method captures the whole target. Since the modes are no longer separated by areas of near-zero probability, the discrete NF methods are now able to accurately capture the target density. In this case, DDS marginally outperforms MFM as measured by the MMD, but this slight improvement in performance comes at the cost of a much higher run-time.
@@ 5.3 Field system
Our first real-world example considers the Allen–Cahn equation, used as a benchmark in [21], and described in Appendix C.4. This fundamental reaction-diffusion equation is central to the study of phase transitions in condensed matter systems. Incorporating random forcing terms or thermal fluctuations allows for a stochastic treatment of the dynamics, capturing the inherent randomness and uncertainties in physical systems. This system leads to a discretized target density of the form
$\log\pi(x)=-\beta\bigg{(}\frac{a}{2\Delta s}\sum_{i=1}^{d+1}(x_{i}-x_{i-1})^{2}+\frac{\Delta s}{4a}\sum_{i=1}^{d}(1-x_{i}^{2})^{2}\bigg{)},$ (18)
where $\Delta s=\frac{1}{d}$, and with boundary conditions $x_{0}=x_{d+1}=0$. In our experiments, we set the dimensionality $d=64$. Meanwhile, parameters values are chosen to ensure bimodality at $x=\pm 1$: $a=0.1$ and inverse temperature $\beta=20$. The bimodality induced by the two global minima complicates mixing when using traditional MCMC updates. Learning the global geometry of the target and using that information to propose transitions facilitates movement between modes. Unlike previous work [e.g., 21], we deliberately choose not to employ an informed base measure. Instead, we opt for a standard Gaussian with no additional information, making the problem significantly more challenging. This choice illustrates the robustness of our approach.
@@
Numerical diagnostics for each method are presented in Table 2. In this case, we use the Kernelized Stein Discrepancy (KSD) as a measure of sample quality [e.g., 43, 23]; see Appendix C.1 for details. While this is not a perfect metric, it does allow us to qualitatively compare the different methods considered.
In this case, the tempering mechanism of our method is crucial for ensuring that the learned flow does not collapse on one of the modes and instead explores both global minima. This is confirmed when plotting the samples generated in the grid in Fig. 3. This experiment demonstrates the capability of our method to capture complex multi-modal densities, even without an informed base measure, at a significantly lower computational cost (e.g., 10-25x faster) than competing methods. It is worth noting that MFM (and the other two benchmarks) significantly outperformed NF-MCMC in this example, despite the similarities between MFM and NF-MCMC. While we tested various hyperparameter configurations for NF-MCMC, we were not able to find a setting that achieved comparbale results.
5.4 Log-Gaussian Cox point process
Bayesian inference for high-dimensional spatial models is known to be challenging. One such model is the log-Gaussian Cox point process, which here we use to model the locations of 126 Scots pine saplings in a natural forest in Finland. See Appendix C.5 for full details. The target space is discretized to an $M=40\times 40$ regular grid, rendering the target dimension $d=1600$. In Table 2, we report diagnostics for each algorithm.
In this case, the lack of multimodality in the target makes it a good fit for non-tempered schemes. Similar to the previous example, NF-MCMC is unable to obtain an accurate approximation to the target distribution. We suspect that this may be a result of non-convergence: due to memory issues, it was not possible to run NF-MCMC (or FAB) for more than $K=10^{3}$ iterations. This also explains the (relatively) smaller run times of these algorithms in this example. By a small margin, DDS provides the best approximation of the target, slightly outperforming MFM and FAB. Meanwhile, MFM provides a good approximation to the target at a lower computational cost with respect to its counterparts.
@@ 5.4 Log-Gaussian Cox point process
FAB $0.14\pm 0.42$ $1.78\pm 0.42$ $3880\pm 7.19$ $1.55$e-1±$\pm 0.06$ $52.3\pm 2.02$ $1040\pm 2.78$
Table 2: Real examples diagnostics where Diagonstic for the two real data examples. KSD U-stat and V-stat are the Kernel Stein Discrepancy U- and V-statistics between the target and samples generated from the learned flow. Results are averaged and empirical 95% confidence intervals over 10 independent runs.
@@ 6 Conclusion
Summary. In this paper, we introduced Markovian Flow Matching, a new approach to sampling from unnormalized probability distributions that augments MCMC with CNFs. Our method combines a local Markov kernel with a non-local, flow-informed Markov kernel, which is adaptively learned during sampling using FM. It also incorporates an adaptive tempering mechanism, which allows for the discovery of multiple target modes. Under mild assumptions, we established convergence of the flow network parameters to a local optimum of the FM objective. Finally, we benchmarked the performance of our algorithm on several examples, illustrating comparable performance to other state-of-the-art methods, often at a fraction of the computational cost.
Limitations. We highlight three limitations of our work. First, our theoretical result established convergence of the flow network parameters obtained via MFM to a local minimum of the FM objective. Further work is required to understand how well these local minima generalize, in order to accurately quantify how accurately the corresponding CNF captures the target posterior. Second, we did not establish non-asymptotic convergence rates for our method. Finally, since it was not the main focus of this work, we did not explore in great detail other choices of architecture for the flow network. We expect that, for certain targets, this could have a significant impact on the performance of MFM. This remains an interesting avenue for future research.
@@ Appendix A Flow-Informed Markov Chain Monte Carlo Methods
Our proof follows closely the proof of [51, Proposition 1]. Let $\theta^{*}$ be a minimiser of the CFM objective in (4), which we recall is given by
@@
We say that $\smash{\hat{\theta}}$ is stability point of this ODE if, given the initial condition $\smash{\theta(0)=\hat{\theta}}$, the ODE admits the unique solution $\smash{\theta(t)=\hat{\theta}}$ for all $t\geq 0$. Naturally, the minimiser $\theta^{*}$ is a stability point of this ODE, since $\smash{\left.\nabla_{\theta}\mathcal{J}(\theta;\pi)\right|_{\theta=\theta^{*}}=0}$. Meanwhile, we call $\Theta$ the domain of attraction of $\theta^{*}$ if, given the initial condition $\theta(0)\in\Theta$, the solution $\theta(t)\in\Theta$ for all $t\geq 0$, and $\theta(t)$ converges to $\theta^{*}$ as $t\rightarrow\infty$.

### v2

@@ 3.2 Adaptive MCMC with Flow Matching
The output of Algorithm 1 is a vector of parameters $\theta_{K}$ which defines a CNF $\smash{(\phi_{t}^{\theta_{K}})_{t\in[0,1]}}$. Under the assumption that the parameter estimate converges, that is, $\smash{\theta_{K}\rightarrow\theta^{*}_{\mathrm{global}}}$ as $K\rightarrow\infty$, where $\smash{\theta^{*}_{\mathrm{global}}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathcal{J}(\theta;\pi)}$ is the global minimizer of the CFM objective $\mathcal{J}(\theta;\pi)$ in (5), this CNF is guaranteed to generate a probability path $(p_{t}^{\theta})_{t\in[0,1]}$ which transports samples from the reference $p_{0}$ to the true target $\pi$ [e.g., 45].
In practice, the objective $\smash{\mathcal{J}(\theta;\pi)}$ is highly non-convex, and thus it is not possible to establish a convergence result of this type without imposing unreasonably strong assumptions on the vector field $\smash{(v_{t}^{\theta})_{t\in[0,1]}}$. This being said, it is reasonable to ask whether $\theta_{K}$ converges to a local optimum of the CFM objective. We now answer this question in the affirmative. In particular, under mild regularity conditions, Proposition 3.1 guarantees that $\theta_{K}\rightarrow\theta^{*}$ almost surely as $K\rightarrow\infty$, where $\theta^{*}$ denotes a local minimum of the CFM objective. This proposition closely mirrors [54, Proposition 1]. Its proof, which relies on a classical result in [6, Theorem 3.17], is provided in Appendix B.
@@ 4 Related work
In recent years, a number of works have proposed algorithms which combine MCMC techniques with NFs; see, e.g., [28, 2] for recent surveys. Broadly speaking, these algorithms fall into two distinct categories. NeutraMCMC methods leverage NFs as reparameterization maps which simplify the geometry of the target distribution, before running (local) MCMC samplers in the latent space. This technique was pioneered in the seminal paper [59], and since been investigated in a number of different works [e.g., 33, 54, 58, 44, 84, 82, 68, 13]. Flow MCMC methods, meanwhile, utilize the pushforward of the base distribution through the NF as an (independent) proposal within an MCMC scheme. This approach was first studied by [3], and further extended in [57, 4, 31].
More recently, [24, 67] have introduced adaptive MCMC schemes which combine local MCMC samplers (e.g., MALA or HMC), with a non-local, flow-informed proposal (IMH or i-SIR); see also [35]. Our algorithm combines aspects of both neutraMCMC and flow MCMC methods and, unlike any existing approach, make use of a CNF (as opposed to a discrete NF), by leveraging the conditional flow matching objective. The use of NFs within other Monte Carlo algorithms has also been the subject of recent interest. For example, [5, 50] consider augmenting SMC with NFs, while [19, 52] use NFs (or diffusion models) within AIS.
Although less directly comparable to our own approach, several other recent works have proposed to use (controlled) diffusion processes to sample from unnormalized probability distributions. Such works include Zhang and Chen [85], who introduce the path integral sampler, Vargas et al. [75], who propose the denoising diffusion sampler, and Zhang et al. [83], who introduce generative flow samplers. Some other relevant contributions in this direction include [73, 8, 63, 76, 75, 77, 74, 78, 60, 16, 1, 69].
@@ 5 Experiments
In this section, we evaluate the performance of MFM (Algorithm 1) on two synthetic and two real data examples. Our method is benchmarked against four relevant methods. The Denoising Diffusion Sampler [DDS; 75] is a VI method which approximates the reversed diffusion process from a reference distribution to an extended target distribution by minimizing the KL divergence. Adaptive Monte Carlo with Normalizing Flows [NF-MCMC; 24] is an augmented MCMC scheme which uses a mixture of MALA and adaptive transition kernels learned using discrete NFs. Flow Annealed Importance Sampling Bootstrap [FAB; 52] is an augmented AIS scheme minimizing the mass-covering $\alpha$-divergence with $\alpha=2$. Finally, Adaptive Tempered SMC (AT-SMC), i.e. the SMC algorithm described in [18] using a MALA transition kernel and a sequence of annealed distributions chosen adaptively by solving (17).
@@
$\text{NN}^{*}(t;\theta_{3})v_{t}^{\theta}(x)=\text{NN}(x,t;\theta_{1})+\text{NN}(t;\theta_{2})\times\nabla\log\pi(x),$ (18)
where the neural networks are standard MLPs with 2 hidden layers, using a Fourier feature augmentation for $t$ [71], and where $\text{NN}^{*}$ outputs a real value that reweights the vector field output using the time component. This architecture is also used by DDS [75, Section 4]. Meanwhile, FAB and NF-MCMC use rational quadratic splines [21]. Flows are trained using Adam [40] with a linear decay schedule terminating at $\varepsilon_{K}=0$. We report results for all methods averaged over 10 independent runs with varying random seeds. Code to reproduce the experiments is provided at https://github.com/albcab/mfm.
@@
Our first example is a mixture of four Gaussians, evenly spaced and equally weighted, in two-dimensional space. The four mixture components have means $(8,8)$, $(-8,8)$, $(8,-8)$, $(-8,-8)$, and all have identity covariance. This ensures that the modes are sufficiently separated to mean that jumping between modes requires trajectories over sets with close to null probability. Given the synthetic nature of the problem, we can measure approximation quality using the Maximum Mean Discrepancy (MMD) [e.g., 29]; see Appendix C.1 for details. We can also include, as a benchmark, the results for an approximation learned using FM with true target samples. Diagnostics for all models are presented in Table 1, and learned flow samples in Figure 1. Further algorithmic details and results are provided in Appendix C.2.
@@
In this experiment, only our method (Figure 1(a)) and DDS (Figure 1(c)) learn the fully separated modes, reflecting the greater expressivity of CNFs in comparison to the discrete NFs used in, e.g., NF-MCMC (Figure 1(d)). It is worth noting that DDS provides a closer approximation to the real target than MFM and, notably, even FM trained using true target samples (top row). Given that both methods use the same network architecture but a different learning objective, this suggests a potential limitation with the FM objective, at least when using this network architecture. This being said, MFM is notably more efficient than DDS (as well as the other methods) in terms of total computation time. While this is not a critical consideration in this synthetic, low-dimensional setting, it is a significant advantage of MFM in higher-dimensional settings involving real data (e.g., Section 5.3 and Section 5.4).
@@
The second experiment is a mixture of bivariate Gaussians with 16 mixture components. This is a modification of the 4-mode example, with contrasting qualities that illustrate other characteristics of each of the presented methods. In this case, the modes are evenly distributed on $[-16,16]^{2}$, with random log-normal variances. The number of modes reduces the size of sets of (near) null probability between the modes, making jumping between them easier. To increase the difficulty of this model, all methods are initialized on a concentrated region of the sampling space. Diagnostics are presented in Table 1 and learned flow samples in Figure 2. Further details are provided in Appendix C.3.
In this example, DDS collapses to the modes closest to the initial positions while our method captures the whole target. Since the modes are no longer separated by areas of near-zero probability, the discrete NF methods are now able to accurately capture the target density. In this case, FAB marginally outperforms MFM as measured by the MMD, but this slight improvement in performance comes at the cost of a much higher run-time.
@@ 5.3 Field system
Our first real-world example considers the stochastic Allen–Cahn model [7], used as a benchmark in [24], and described in Appendix C.5. This fundamental reaction-diffusion equation is central to the study of phase transitions in condensed matter systems. Incorporating random forcing terms or thermal fluctuations allows for a stochastic treatment of the dynamics, capturing the inherent randomness and uncertainties in physical systems. This model leads to a discretized target density which takes the form
$\log\pi(x)=-\beta\bigg{(}\frac{a}{2\Delta s}\sum_{i=1}^{d+1}(x_{i}-x_{i-1})^{2}+\frac{b\Delta s}{4}\sum_{i=1}^{d}(1-x_{i}^{2})^{2}\bigg{)},$ (19)
with $\Delta s=\frac{1}{d}$, and boundary conditions $x_{0}=x_{d+1}=0$. In our experiments, we take $d=64$. Meanwhile, following [24], other parameter values are chosen to ensure bimodality at $x=\pm 1$: $a=0.1$, $b=1/a=10$, and $\beta=20$. The bimodality induced by the two global minima complicates mixing when using traditional MCMC updates. Learning the global geometry of the target and using that information to propose transitions facilitates movement between modes. Unlike previous work [e.g., 24], we deliberately choose not to employ an informed base measure. Instead, we opt for a standard Gaussian with no additional information, making the problem significantly more challenging. This choice illustrates the robustness of our approach.
Numerical diagnostics for each method are presented in Table 2. In this case, we use the Kernelized Stein Discrepancy (KSD) as a measure of sample quality [e.g., 46, 26]; see Appendix C.1 for details. While this is not a perfect metric, it does allow us to qualitatively compare the different methods considered.
In this case, the tempering mechanism of our method is crucial for ensuring that the learned flow does not collapse on one of the modes and instead explores both global minima. This is confirmed when plotting the samples generated in the grid in Figure 3. This experiment demonstrates the ability of our method to capture complex multi-modal densities, even without an informed base measure, at a significantly lower computational cost (e.g., 10-25x faster) than competing methods. Indeed, while FAB was the best performing method in this experiment as measured by the KSD, it failed to capture both of the modes in the target distribution, and required a much greater total computation time (see Table 2).
It is worth noting that MFM (and the other two benchmarks, DDS and FAB) significantly outperformed NF-MCMC in this example, despite the similarities between MFM and NF-MCMC. While we tested various hyperparameter configurations for NF-MCMC, we were not able to find a setting that achieved comparable results in the absence of an informed base measure.
@@
5.4 Log-Gaussian Cox process
Bayesian inference for high-dimensional spatial models is known to be challenging. One such model is the log-Gaussian Cox process (LGCP) introduced in [53], which is used to model the locations of 126 Scots pine saplings in a natural forest in Finland. See Appendix C.6 for full details. The target space is discretized to a $M=40\times 40$ regular grid, rendering the target dimension $d=1600$. In Table 2, we report diagnostics for each algorithm.
In this case, the lack of multimodality in the target makes it a good fit for non-tempered schemes. Similar to the previous example, NF-MCMC is unable to obtain an accurate approximation to the target distribution. We suspect that this may be a result of non-convergence: due to memory issues, it was not possible to run NF-MCMC (or FAB) for more than $K=10^{3}$ iterations. This also explains the (relatively) smaller run times of these algorithms in this example. By a small margin, DDS provides the best approximation of the target, slightly outperforming MFM and FAB. Meanwhile, MFM provides a good approximation to the target at a lower computational cost with respect to its competitors.
@@ 5.4 Log-Gaussian Cox process
FAB $0.14\pm 0.42$ $1.78\pm 0.42$ $3880\pm 7.19$ $1.55$e-1±.$\pm.06$ $52.3\pm 2.02$ $1040\pm 2.78$
AT-SMC $1.61\pm 2.33$ $18.4\pm 2.35$ $4.13\pm 0.30$ $1.39$e-2±.$\pm.01$ $25.0\pm 0.12$ $6.11\pm 0.44$
Table 2: Diagnostics for the two real data examples. KSD U-stat and V-stat are the Kernel Stein Discrepancy U- and V-statistics between the target and samples generated from the learned flow. Results are averaged and empirical 95% confidence intervals over 10 independent runs.
@@ 6 Conclusion
Summary. In this paper, we introduced Markovian Flow Matching, a new approach to sampling from unnormalized probability distributions that augments MCMC with CNFs. Our method combines a local Markov kernel with a non-local, flow-informed Markov kernel, which is adaptively learned during sampling using FM. It also incorporates an adaptive tempering mechanism, which allows for the discovery of multiple target modes. Under mild assumptions, we established convergence of the flow network parameters output by our algorithm to a local optimum of the FM objective. We also benchmarked the performance of our algorithm on several examples, illustrating comparable performance to other state-of-the-art methods, often at a fraction of the computational cost.
Limitations and Future Work. We highlight three limitations of our work. First, our theoretical result established convergence of the flow network parameters obtained via MFM to a local minimum of the FM objective. Further work is required to understand how well these local minima generalize, in order to accurately quantify how accurately the corresponding CNF captures the target posterior. Second, we did not establish non-asymptotic convergence rates for our method. Finally, since it was not the main focus of this work, we did not explore in great detail other choices of architecture for the flow network. We expect that, for certain targets, this could have a significant impact on the performance of MFM. Indeed, a promising avenue for further research lies in developing tailored CNFs designed for particular posterior distributions. This approach would go beyond the current practice of including the gradient of the log-posterior and instead exploit unique characteristics intrinsic to each model when constructing the flow.
@@ Appendix A Flow-Informed Markov Chain Monte Carlo Methods
Our proof follows closely the proof of [54, Proposition 1]. Let $\theta^{*}$ be a minimizer of the CFM objective in (4), which we recall is given by
@@
We say that $\smash{\hat{\theta}}$ is stability point of this ODE if, given the initial condition $\smash{\theta(0)=\hat{\theta}}$, the ODE admits the unique solution $\smash{\theta(t)=\hat{\theta}}$ for all $t\geq 0$. Naturally, the minimizer $\theta^{*}$ is a stability point of this ODE, since $\smash{\left.\nabla_{\theta}\mathcal{J}(\theta;\pi)\right|_{\theta=\theta^{*}}=0}$. Meanwhile, we call $\Theta$ the domain of attraction of $\theta^{*}$ if, given the initial condition $\theta(0)\in\Theta$, the solution $\theta(t)\in\Theta$ for all $t\geq 0$, and $\theta(t)$ converges to $\theta^{*}$ as $t\rightarrow\infty$.